You are going to make a box out of cardboard. The box needs to measure 10 cm by 5 cm by 5 cm; 10% of the final product will be wasted in the cutting process. How much cardboard do you need?


My Answer:

125 cm.
140 cm.
200 cm.
250 cm.
275 cm.

Answer by checkley77(12844)

The original piece of cardboard was:
5+10+5=20
5+5+5=15
20*15=300 cm^2 is the original size of the cardboard.
300-100=200 is the area of the remaining cardboard.